OpenStudy (anonymous):

How many solutions does the following equations have? 1. –7x² + 6x + 3 = 0 2. –8x² – 8x – 2 = 0 Thanks in advance to all answers! :D

OpenStudy (king):

2 each

OpenStudy (hoblos):

1) 2 solutions 2) a double solution

OpenStudy (experimentx):

both equations are quadradic equations ... so at most they have two solutions, either both real or both imaginary

OpenStudy (king):

both are quadratic equation and therefor cannot have more than 2 solutions........

OpenStudy (anonymous):

both are quadratic equation and therefor cannot have more than 2 solutions........ so 2

OpenStudy (experimentx):

in the latter case, it has only one solution .. .i.e two roots coincide with each other

OpenStudy (anonymous):

for (1) , b^2 -4ac = 36-4(-7)(3) >0 therefore, 2 solutions for (2), b^2 -4ac =64-4(-8)(-2) =0 therefore, 1 repeated root